Dear tutor,
Please help with this
2/5x=2/7
solve for x
I figured out the answer = 5/7
but how do I show the work?

Answer by unlockmath(1688)

Hello,
Looks like you have it good!
Let's first get this so there are no fractions. So we need to multiply both sides. I have only 1 question. Is the x in like this:
2/5x=2/7 or 2/5(x)=2/7 I'll solve both. The first one is multiplied by 35x:
(35x)2/5x=35x(2/7) This will turn out to be:
14=10x Divide both sides by 10 will give us:
x=14/10 or 7/5
2/5(x)=2/7 This we'll multiply by 35
(35)2/5(x)=35(2/7) This will turn out as:
14x=10 Divide both sides by 14 will give us:
x=10/14 or 5/7
Hope this helps you out. Let me know.
RJ Toftness
www.math-unlock.com